# OVA-Det: Open Vocabulary Aerial Object Detection with Image-Text Collaboration

### v2

@@ 5 Conclusion
In this paper, we propose OVA-Det, aiming to develop an efficient open-vocabulary detector for aerial object detection. To this end, we focus on exploring image-text collaboration. Specifically, we incorporate class semantics into the detector and introduce a image-to-text alignment loss to enhance image-text alignment, thereby eliminating the predefined category constraints of traditional detectors. We further propose TG-FE and TG-QE, forming a text-guided encoder-decoder framework. This architecture makes the model focus on class-relevant features, allowing it to utilize rich textual information for improved detection performance. Moreover, by avoiding computationally intensive modules, OVA-Det inherits the efficiency of its baseline, meets the real time demand of aerial detection, and is therefore more suitable for open scenarios. In future work, we will further explore text guidance, extending word-level cues to phrases and even sentences to provide more detailed guidance information, thereby enhancing open detection capabilities while broadening the model’s applicability.
Supplementary Material
@@
6 Dataset Details
6.1 Zero-shot Detection Dataset Details
To ensure a fair experimental comparison, we followed the state-of-the-art DescReg [40] for open aerial object detection to organize our experiments. We evaluate OVA-Det zero-shot detection performance on three benchmark datasets. The dataset details as follows:
@@ 6.1 Zero-shot Detection Dataset Details
6.1.1 Dataset Crop Details
@@
6.1.2 Dataset Base/Novel Split Details
@@ 7 Qualitative Results
As shown in Fig 6, we present qualitative results of OVA-Det on the xView, DIOR, and DOTA datasets. The images in the first, second, and third rows correspond to xView, DIOR, and DOTA, respectively. These results demonstrate that our proposed method breaks the category limitations of traditional detectors and can accurately detect novel classes. As can be seen, even for targets with inconspicuous visual features, such as the bus and motorboat in the first row, our proposed OVA-Det can still detect them.
8 Implementation Details
8.1 Training Details
@@
The OVA-Det implementation is based on RT-DETR [46] with a ResNet50 [13] backbone, using the default hyperparameters provided by RT-DETR. In the query selection module, we replace classification scores with image-text similarity scores to select the top 500 encoder features, which are then used to initialize the object queries for the decoder. We also use these similarity scores instead of classification scores for label assignment. We employ a one-stage training scheme, excluding novel classes from the training sets of all three datasets for training. The model is evaluated on both GZSD and ZSD tasks using the test sets of these three datasets.
@@
Following the approach in [22], we adopt a random class sampling procedure during training. Specifically, we shuffle the order of input text each time, preventing the model from relying on fixed positions of class embeddings. We first sample the positive classes that appear in the images, and then randomly sample the remaining negative classes to reach the required number of training classes. If there are insufficient negative classes, we use an empty token as a placeholder and mask the corresponding class embeddings in subsequent operations.
8.2 Details of the compared method
@@
Due to the unavailability of the code from prior work [40], we directly use the results reported in their publication for experiments under their dataset split. YOLO-World [4] is the most effective OVD method for natural images. We compare our model against it to demonstrate the OVA-Det’s suitability for aerial OVD tasks. Specifically, we follow the training procedures detailed in their published paper and available code, conducting experiments under settings similar to ours.
To highlight the advantages of our approach, we use the versions of each method that have been pretrained on the largest amount of data (for example, a YOLO-World model pretrained on Objects365V1 [31], GQA [16], Flickr [27], and CC3M [32]). We then fine-tune these models on the three aerial image datasets.
9 More Experimental Results
9.1 Comparison with RT-DETR
As described in the main text, our method follows the RT-DETR framework. Therefore, we compare OVA-Det and RT-DETR in terms of model size, detection accuracy, and inference speed. Since RT-DETR cannot be applied to open vocabulary scenarios, we conducted closed-world comparison experiments on Visdrone. The corresponding results are listed in Table 7.
Based on the experimental results, our method can also improve detection accuracy in closed scenarios. To accommodate open scenarios, additional parameters and computational overhead were introduced. Image-to-text alignment loss introduced about 2M parameters, TG-FE and TG-QE introduced about 6M parameters. However, since the additional modules are all linear structures, the hardware inference speed is relatively fast. Therefore, the model’s latency increase is minimal, with an FPS reduction of about 5 frames. Ultimately, it maintains an inference speed of 36 FPS, which is sufficient to meet the real-time inference requirements of typical videos (25–30 FPS).
9.2 Performance on UAVDT dataset
UAVDT [9] is also a benchmark for evaluation in aerial object detection. The dataset comprises 23,258 training images and 15,069 testing images with a resolution of $1024\times 540$, covering 3 classes.
Besides Visdrone, to further demonstrate the performance of OVA-Det on traditional aerial detection tasks, we evaluate our method on the UAVDT dataset. As shown in Table 8, OVA-Det achieves the highest mAP performance, surpassing the best-compared method by 0.5% in mAP. In summary, our proposed OVA-Det not only adapts to open-vocabulary aerial object detection scenarios but also retains excellent performance in closed settings.
Method Params mAP FPSbs=1
RT-DETR(ResNet-50) 41M 58.9 41
OVA-Det(ResNet-50) 49M 59.7 36
Table 7: Comparison of OVA-Det and RT-DETR on the Visdrone Dataset.
Method BackBone Resolution mAP
DMNet[19] ResNet-50 1000 $\times$ 600 24.6
ClusDet[38] ResNet-50 1000 $\times$ 600 26.5
GLSAN[6] ResNet-50 1000 $\times$ 600 28.1
ARMNet[33] ResNet-50 1500 $\times$ 800 30.4
RT-DETR[46] ResNet-50 800 $\times$ 800 39.5
UFPMP-Det[15] ResNet-50 1000 $\times$ 600 38.7
GFL V1(CEASC)[8] ResNet-18 1024 $\times$ 540 30.9
YOLO-World[4] YOLOv8-L 800 $\times$ 800 35.8
OVA-Det ResNet-50 800 $\times$ 800 41.0
Table 8: Comparison with the state-of-the-art on the UAVDT dataset.

### v3

@@ V Conclusion
In this paper, we propose RT-OVAD, aiming to develop an efficient open-vocabulary detector for aerial object detection. To this end, we focus on exploring image-text collaboration. Specifically, we incorporate class semantics into the detector and introduce a image-to-text alignment loss to enhance image-text alignment, thereby eliminating the predefined category constraints of traditional detectors. We further propose TG-FE, VG-TR and TG-QE, forming a image-text collaboration encoder-decoder framework. This architecture makes the model focus on class-relevant features, allowing it to utilize rich textual information for improved detection performance. Moreover, by avoiding computationally intensive modules, RT-OVAD inherits the efficiency of its baseline, meets the real time demand of aerial detection, and is therefore more suitable for open scenarios. In future work, we will further explore text guidance, extending word-level cues to phrases and even sentences to provide more detailed guidance information, thereby enhancing open detection capabilities while broadening the model’s applicability.
@@
Appendix A Dataset Details
A-A Zero-shot Detection Dataset Details
To ensure a fair experimental comparison, we followed the state-of-the-art DescReg [18] for open aerial object detection to organize our experiments. We evaluate RT-OVAD zero-shot detection performance on three benchmark datasets. The dataset details as follows:
@@ A-A Zero-shot Detection Dataset Details
A-A1 Dataset Crop Details
@@
A-A2 Dataset Base/Novel Split Details
@@ Appendix B Qualitative Results
As shown in Fig 6, we present qualitative results of RT-OVAD on the xView, DIOR, and DOTA datasets. The images in the first, second, and third rows correspond to xView, DIOR, and DOTA, respectively. These results demonstrate that our proposed method breaks the category limitations of traditional detectors and can accurately detect novel classes. As can be seen, even for targets with inconspicuous visual features, such as the bus and motorboat in the first row, our proposed RT-OVAD can still detect them.
Appendix C Implementation Details
C-A Training Details
@@
The RT-OVAD implementation is based on RT-DETR [25] with a ResNet50 [41] backbone, using the default hyperparameters provided by RT-DETR. In the query selection module, we replace classification scores with image-text similarity scores to select the top 500 encoder features, which are then used to initialize the object queries for the decoder. We also use these similarity scores instead of classification scores for label assignment. For open-vocabulary aerial detection, following the protocol in [21], all models are first pretrained on the large-scale LAE-1M dataset and subsequently evaluated on the DIOR, DOTAv2.0, and xView benchmarks. In the zero-shot setting, we adopt a one-stage training strategy, where novel categories are excluded from the training sets of all three datasets. The models are evaluated on both generalized zero-shot detection (GZSD) and zero-shot detection (ZSD) tasks using the validation sets of the corresponding benchmarks.
@@
Following the approach in [17, 16], we adopt a random class sampling procedure during training. Specifically, we shuffle the order of input text each time, preventing the model from relying on fixed positions of class embeddings. We first sample the positive classes that appear in the images, and then randomly sample the remaining negative classes to reach the required number of training classes. If there are insufficient negative classes, we use an empty token as a placeholder and mask the corresponding class embeddings in subsequent operations.
C-B Details of the compared method
For open-vocabulary aerial detection, we compare RT-OVAD with four representative open-vocabulary detectors, including both general-purpose [17, 16, 15] and aerial image–tailored models [21]. Except for RT-OVAD and YOLO-World, all results are directly taken from the LAE-DINO [21] paper for consistency and fairness.
@@ C-B Details of the compared method
For zero-shot aerial detection, Due to the unavailability of the code from prior work [18], we directly use the results reported in their publication for experiments under their dataset split. We also compare our model against the YOLO-World [16], Grounding DINO [15] and LAE-DINO [21] to demonstrate the RT-OVAD’s suitability for aerial zero-shot detection tasks. Specifically, we follow the training procedures detailed in their published paper and available code, conducting experiments under settings similar to ours.
C-C Limitation Analysis
While RT-OVAD achieves strong performance across three prevalent benchmark tasks, several limitations persist: (1) Limited textual annotations. Current open-vocabulary aerial benchmarks provide only category-level labels. The lack of phrase- or sentence-level descriptions limits the richness and expressiveness of textual guidance during both training and inference. (2) Potential instability during zero-shot evaluation. Under zero-shot settings, we observe that the model exhibits mild performance fluctuations. We attribute these variations to: (i) the sparsity of textual prompts for novel classes, (ii) the long-tailed distribution common in aerial categories, and (iii) inherent stochasticity in the training and inference processes. To improve evaluation stability, we report the average results over three independent runs. (3) High recall but low precision on novel classes. In certain base/novel class split settings, we observe a substantial number of false positives, where background regions are mistakenly classified as novel objects. This may indicate that the image-to-text alignment loss lacks explicit constraints for background regions, making it difficult for the model to accurately distinguish novel targets from complex background features.
Despite these limitations, RT-OVAD still delivers state-of-the-art accuracy while maintaining real-time inference speed. Future work may consider: (i) enriching benchmark datasets with fine-grained textual annotations, and (ii) exploring prompt-learning strategies or background-aware loss functions to improve model discrimination.